OpenStudy (anonymous):

Fourteen more than three times a number is 2

OpenStudy (anonymous):

3x+14=2

OpenStudy (uri):

14 +3x=2 3x=2-14 3x=-12 x=-12/3 x=-4

OpenStudy (uri):

If you want to check if your answer is right,plug -4 in the equation: 14+3(-4)=2 14-12=2 @Maydelisx3